A PROOF OF ATIYAH’S CONJECTURE ON CONFIGURATIONS OF FOUR POINTS IN EUCLIDEAN THREE-SPACE

MICHAEL EASTWOOD PAUL NORBURY

Abstract. From a configuration of n points in Euclidean three-space, Atiyah constructed a determinant and conjectured that it was always non-zero. In this article we prove the conjecture for the case of four points.